# GAL-MAD: Towards Explainable Anomaly Detection in Microservice Applications Using Graph Attention Networks

### v1

@@ Abstract
The transition to microservices has revolutionized software architectures, offering enhanced scalability and modularity. However, the distributed and dynamic nature of microservices introduces complexities in ensuring system reliability, making anomaly detection crucial for maintaining performance and functionality. Anomalies stemming from network and performance issues must be swiftly identified and addressed. Existing anomaly detection techniques often rely on statistical models or machine learning methods that struggle with the high-dimensional, interdependent data inherent in microservice applications. Current techniques and available datasets predominantly focus on system traces and logs, limiting their ability to support advanced detection models. This paper addresses these gaps by introducing the RS-Anomic dataset generated using the open-source RobotShop microservice application. The dataset captures multivariate performance metrics and response times under normal and anomalous conditions, encompassing ten types of anomalies. We propose a novel anomaly detection model called Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD), leveraging Graph Attention and Long Short-Term Memory architectures to capture spatial and temporal dependencies in microservices. We utilize SHAP values to localize anomalous services and identify root causes to enhance explainability. Experimental results demonstrate that GAL-MAD outperforms state-of-the-art models on the RS-Anomic dataset, achieving higher accuracy and recall across varying anomaly rates. The explanations provide actionable insights into service anomalies, which benefits system administrators.
@@ 1 Introduction
The migration from monolithic to microservice architecture has become a prevalent trend in cloud applications due to benefits such as improved scalability, modularity, and autonomous components that simplify development. A few enterprise companies using microservice architecture include Amazon, Netflix, and Uber. A microservice application consists of loosely coupled, independently deployable services, with each microservice performing a specific functionality or a business objective. Typically the performance of microservices are gaurded by Service Level Obejctives (SLOs).
@@
In this research, we present the RS-Anomic dataset, which contains multivariate data from a microservice application and a framework for detecting anomalous behaviour of a microservice application and localizing the affected microservice. We utilized RobotShop Instana (2022) as the microservice application for this dataset. RobotShop is a small-scale microservice application that aims to teach containerized application orchestration and monitoring methods. RS-Anomic provides performance metrics on memory, CPU utilization, file Input/Output, and network as a multivariate time series for each service. These metrics can be valuable in developing novel anomaly detection methods that explore inter-dependencies between services and their performance metrics.
Moreover, we introduce a new model called Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) for detecting anomalies in microservices. This model is an encoder-decoder architecture trained using the performance metrics of the microservice system during normal operation. When provided with the performance metrics over a specific period, the model learns to reconstruct this input accurately. During inference, if the system’s performance metrics indicate anomalous behaviour, the model exhibits a higher reconstruction loss, aiding in identifying such anomalies. The encoder component employs Graph Attention layers to capture the dependencies among cooperating microservices, and Long Short-Term Memory layers to recognize the temporal patterns in the operations of the microservice system. The decoder mirrors the structure of the encoder. To assess the performance of our model, we compare it with state-of-the-art anomaly detection models using the RS-Anomic dataset, highlighting the importance of capturing interdependencies within the microservices domain. Additionally, we utilize SHapley Additive exPlanations Lundberg and Lee (2017a) for root cause analysis, which reinforces the model’s predictions and enhances its practical utility.
@@ 2.1 Datasets for Anomaly Detection
Performance and response time features are crucial in detecting and localizing anomalies in microservices. Large datasets are available for anomaly detection using microservice trace logs Qiu et al. (2020); Huye et al. (2023); Lee et al. (2024). By analyzing trace logs, slow or erroneous services can be identified. Traces serve as a unified response time metric for service calls, encompassing both latency and performance aspects. However, root cause analysis using traces are often limited to identifying the specific service call and the anomalous service. To analyze system level resource usage multivariate data consisting of performance metrics are required. Response time is a key component of latency related SLOs and is essential for anomaly detection. NAB dataset collection from Numeta Lavin and Ahmad (2015) consists of a dataset built using AWS server metrics collected by the Amazon CloudWatch service. However, this dataset lacks response time related metrics. In contrast, a more comprehensive dataset combining system resource utilization and response time was introduced by Luo et al. Luo et al. (2021), however it does not include network resource usage data. Since detecting fine network anomalies requires detailed network resource usage metrics, future datasets should incorporate these to enhance anomaly detection capabilities.
@@ 2.2 Anomaly Detection
In early anomaly detection techniques, statistical methods were commonly used. Researchers analyzed the probability distribution of data points, identifying points outside the defined range of normal behaviour as outliers, and therefore, anomalies. Traditional techniques mainly include non-deep learning and log analysis techniques. Log analysis is complicated as logs are unstructured for large systems and challenging to perform in real time. In addition, to log analysis techniques, traditional machine learning methods have been used alongside deep learning approaches.
Recent research has shown that deep learning-based approaches generally outperform traditional and statistical methods in anomaly detection. However, anomaly detection remains a difficult task due to the lack of anomalous data and the highly imbalanced nature of available datasets, as anomalies are rare compared to normal behaviour. To overcome this issue, unsupervised approaches were utilized to distinguish abnormal from normal patterns Liu et al. (2020); Nguyen et al. (2021); Chen et al. (2017).
There are several advanced models for detecting anomalies, including the Graph Deviation Network (GDN) Deng and Hooi (2021), the Multivariate Anomaly Detection with Generative Adversarial Network (MAD-GAN)Li et al. (2019), and KitsuneMirsky et al. (2018). In Chen et al. (2017) introduced autoencoder ensembles for unsupervised outlier detection, demonstrating that neural network-based anomaly detection techniques yield high accuracy. Their model reconstructs normal data, identifying anomalies as instances that cannot be accurately reconstructed. The problem of accurately learning complex dependencies in multivariate data is addressed by Graph Attention Networks (GAT) Velickovic et al. (2017). GAT is further extended in GDN Deng and Hooi (2021), using a cosine similarity to create a graph structure to represent the pairwise relationship between data points. It then uses graph attention-based forecasting to predict future values and calculates the graph deviation score to evaluate its performance. MAD-GAN Li et al. (2019) is designed to detect anomalies in time series data produced by Cyber-Physical Systems. This model is trained using normal data and uses a Long-Short Term Memory-Recurrent Neural Network discriminator and a reconstruction-based method to calculate the anomaly score for each sample. Kitsune Mirsky et al. (2018) is an unsupervised Network Intrusion Detection System that can detect local network attacks without supervision. Kitsune uses a collection of variational autoencoders trained to replicate network traffic patterns and gradually improve their performance with Kitsune’s core algorithm, KitNET, through incremental training and efficient pattern tracking.
@@ 2.3 Root Cause Analysis
There is a lack of research exploring the use of explainability in root cause analysis for anomaly detection. By employing a graph-based anomaly detection method that preserves the microservice architecture, we investigate the potential of anomaly localization through explainability. We propose locating anomalous services using explainability techniques Lundberg and Lee (2017b) in conjunction with the output of the model. The tension between accuracy and interpretability in complex models has led to the development of methods for interpreting predictions. Lundberg et. al. introduced SHAP, a unified framework that assigns importance values to input features in a prediction. This framework offers a unique solution with desirable properties, enhancing existing methods to provide improved alignment with human intuition.
@@ 3.1.1 Testbed and Data Collection Environment
The RS-Anomic dataset was created using RobotShop Instana (2022), a microservice application that implements containerized orchestration and monitoring techniques. RobotShop is an open-source project with an e-commerce web application built using a microservices architecture. All services are deployed as Docker containers on a single server, communicating through a Docker bridge network. The application comprises 12 services, with their dependencies illustrated in Figure 1. This application was used in Harlicaj (2021) to evaluate their model. The testbed was run on a server with a 40-core CPU and 64 GB Memory running Ubuntu 20.04.5. We used Prometheus Rabenstein and Volz (2015) for data acquisition. The Prometheus server was configured to poll data from the microservices every 5 seconds. We instrumented the response times in microservice calls using Prometheus client libraries. The instrumented services are highlighted in Figure 1. Using Prometheus client libraries response time metrics were exposed from the microservices over HTTP endpoints. The cumulative summation of response time and the number of service calls that occurred between data polling for each communication link were recorded. Container Advisor (cAdvisor) Google (2022) was utilized to gather runtime metrics. This tool offers measurements related to resource usage, such as CPU, memory, disk Input/Output (I/O), and network I/O of Docker containers. Additionally, data from cAdvisor was also polled from the Prometheus server. Data collection spanned over six days under normal conditions. For each type of anomalous behaviour, data was collected over 90 minutes by injecting faults into specific services to simulate anomalies. Table 1 shows the features available in the RS-Anomic dataset. To ensure RS-Anomic covers an actual e-commerce application behaviour, we simulated a time-varying load with more users in peak hours and fewer in off-peak hours. We used Locust Locust (2022), an open-source load testing tool, for load generation. The load generation can be done by creating test scenarios that simulate user behaviour and defining the number of virtual users and requests per second that should be generated during the test.
@@ 3.1.2 Anomalous Behaviours Captured in RS-Anomic
We observed that network packet loss, out-of-order packets and high latency anomalies do not clearly distinguish between normal and anomalous data. Therefore, we increased the strength of these anomalies. Figure 2 shows t-SNE plots of comparison between two strengths of anomalous behaviour. The t-SNE plot for 1000ms latency shows clear distinctions between normal and anomalous behaviour for increased strength, whereas 200ms latency anomaly is more challenging to distinguish from normal data points.
(a) 200ms latency
(b) 1000ms latency
Figure 2: t-SNE plots for normal(blue) and anomalous data (orange) of varying strengths.
3.2 GAL-MAD Model
In recent work, we observed that unsupervised approaches have shown promising improvements due to the lack of anomalous data availability. Reproducing anomalies for various microservice applications is often impractical. Therefore, we propose an unsupervised approach, that leverages the normal behaviour performance metrics of services to train the model.
Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) is an autoencoder model Figure 3 illustrates the overall architecture of GAL-MAD. Figure 4 depicts the encoder component consisting of two GAT layers to extract latent information regarding the input feature dependencies and reduce the dimension along the feature axis. The GAT considers attention to neighboring services and features with the help of the adjacency matrix that represents the architecture of the microservice application, Figure 5 visualizes the attention computation for the input features. A many-to-one LSTM is used to reduce the dimension along the time axis. Each LSTM uses a sequence length of 24, translating to 2-minute time windows polled every 5 seconds. A smaller time window would improve responsiveness at the expense of accuracy.

### v2

@@ Abstract
The distributed and dynamic nature of microservices introduces complexities in ensuring system reliability, making anomaly detection crucial. Existing anomaly detection techniques often rely on statistical models or machine learning methods that struggle with high-dimensional and interdependent data. Available datasets have limitations in capturing network and performance metrics of microservices. This work introduces the RS-Anomic dataset generated using the open-source RobotShop microservice application. The dataset captures multivariate performance metrics and response times under normal and anomalous conditions, encompassing ten types of anomalies. We propose a novel anomaly detection model called Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD), leveraging Graph Attention and Long Short-Term Memory architectures to capture spatial and temporal dependencies in microservices. Using SHAP, we explore anomaly localization to enhance explainability. Experimental results demonstrate that GAL-MAD outperforms state-of-the-art models on the RS-Anomic dataset, achieving higher accuracy and recall across varying anomaly rates.
@@ 1 Introduction
The migration from monolithic to microservice architecture has become a prevalent trend in cloud applications due to benefits such as improved scalability, modularity, and autonomous components that simplify development. A few enterprise companies using microservice architecture include Amazon, Netflix, and Uber. A microservice application consists of loosely coupled, independently deployable services, with each microservice performing a specific functionality or a business objective. Typically the performance of microservices are guarded by Service Level Objectives (SLOs).
@@
This research presents the RS-Anomic dataset, which contains multivariate data from a microservice application and a framework for detecting anomalous behaviour of a microservice application and localizing the affected microservice. We utilized RobotShop Instana (2022) as the microservice application for this dataset. RobotShop is a small-scale microservice application that aims to teach containerized application orchestration and monitoring methods. RS-Anomic provides performance metrics on memory, CPU utilization, file Input/Output, and network as a multivariate time series for each service. These metrics can be valuable in developing novel anomaly detection methods that explore inter-dependencies between services and their performance metrics.
Moreover, we introduce a new model called Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) for detecting anomalies in microservices. This model is an encoder-decoder architecture trained using the performance metrics of the microservice system during normal operation. When provided with the performance metrics over a specific period, the model learns to reconstruct this input accurately. During inference, if the system’s performance metrics indicate anomalous behaviour, the model exhibits a higher reconstruction loss, aiding in identifying such anomalies. The encoder component employs Graph Attention layers to capture the dependencies among cooperating microservices, and Long Short-Term Memory layers to recognize the temporal patterns in the operations of the microservice system. The decoder mirrors the structure of the encoder. To assess the performance of our model, we compare it with state-of-the-art anomaly detection models using the RS-Anomic dataset, highlighting the importance of capturing interdependencies within the microservices domain. Additionally, we utilize SHapley Additive exPlanations Lundberg and Lee (2017a) for root cause analysis, reinforcing the model’s predictions and enhancing its practical utility.
@@ 2.1 Datasets for Anomaly Detection
Performance and response time features play a critical role in detecting and localizing anomalies in microservices. Several large-scale datasets are available for anomaly detection using microservice trace logs Qiu et al. (2020); Huye et al. (2023); Lee et al. (2024). These trace logs enable the identification of slow or faulty services by capturing end-to-end service call behaviors. As traces provide unified measurements of response time, including latency and overall performance, they are effective for service-level anomaly detection. However, their utility for root cause analysis is often limited to pinpointing the anomalous service or service call, without offering deeper insights into system-wide issues.
To gain a more holistic understanding of system behavior, multivariate datasets capturing system-level performance metrics are essential. Response time remains a key component of latency-related Service Level Objectives (SLOs) and is central to many anomaly detection techniques. For instance, the NAB dataset collection from Numenta Lavin and Ahmad (2015) includes server metrics collected via Amazon CloudWatch but lacks response time metrics. In contrast, Luo et al. Luo et al. (2021) introduced a more comprehensive dataset that combines system resource utilization with response time data. However, it does not capture network resource usage, which is crucial for detecting subtle or fine-grained network anomalies. Therefore, this research introduces a new multivariate dataset capturing performance metrics of a microservice network to improve anomaly detection effectiveness.
@@ 2.2 Anomaly Detection
Anomaly detection focuses on identifying observations that significantly diverge from normal patterns, a task central to fraud detection, cybersecurity, and healthcare monitoring. Traditional approaches rooted in statistical tests and shallow Machine Learning (ML) classifiers often struggle with high dimensional or unstructured data and the severe class imbalance inherent to anomaly detection tasks. Deep learning has revolutionized anomaly detection in recent years by learning hierarchical representations that capture complex data distributions, yielding more accurate and scalable solutions Pang et al. (2021). A prevalent paradigm employs pre-trained neural networks, such as CNNS or RNNS, as feature extractors, whose embeddings are subsequently fed into ML classifiers like One-Class Support Vector Machines to detect outliers Perera and Patel (2019). Reconstruction based models, including autoencoders and variational autoencoders, train solely on normal data and flag inputs with high reconstruction error as anomalies, Chen et al. (2017); Mirsky et al. (2018). Prediction based techniques leverage forecasting networks often LSTM autoencoders that predict future time steps, denoting large prediction deviations as anomalies Li et al. (2019); Ergen and Kozat (2019). Generative adversarial frameworks learn the distribution of normal data via adversarial training, using both generator reconstruction losses and discriminator confidences for anomaly scoring Li et al. (2019); Schlegl et al. (2019). End-to-end one-class neural methods such as Deep SVDD and its semi-supervised extension DeepSAD unify representation learning and scoring by optimizing hypersphere or ranking objectives to tightly enclose normal instances in latent space Chi and Mao (2024). This research explores the approaches that learn feature representations of normality for anomaly detection, specifically focusing on those based on graph data.
The Graph Deviation Network (GDN) employs a graph neural network architecture to learn inter-variable relationships in multivariate time series via attention-based forecasting, using a structure learned from cosine similarity to compute deviation scores for anomaly detection tasks with both high accuracy and explainability Deng and Hooi (2021). The Multivariate Anomaly Detection with Generative Adversarial Network (MAD-GAN) Li et al. (2019) integrates LSTM-based generator and discriminator networks within a GAN framework to jointly capture temporal dependencies and distributional interactions among variables, producing a novel DR-score that combines reconstruction error and discriminator confidence to flag anomalies. KitsuneMirsky et al. (2018), an unsupervised network intrusion detection system, utilizes the KitNET autoencoder ensemble, trained incrementally on real-time network traffic features extracted by the AfterImage framework, to reconstruct normal traffic patterns, identifying deviations through elevated reconstruction errors and ensemble based thresholding, all with minimal computational overhead suitable for edge devices. Chen et al. Chen et al. (2017) introduced autoencoder ensembles for unsupervised outlier detection, demonstrating that randomly varying the connectivity architectures of autoencoders and employing adaptive sampling enhances robustness to noise and improves detection performance across benchmark datasets. Underpinning these graph-based models, Graph Attention Networks (GAT) extend traditional graph neural networks by incorporating masked self-attention mechanisms that assign learnable importance weights to neighboring nodes, enabling inductive learning on graphs with complex, heterogeneous structures without prior knowledge of the graph topology Velickovic et al. (2018).
A growing body of work addresses anomaly detection in microservice systems by leveraging unstructured logs, structured performance metrics, or a combination of both. Log-based methods evolved from statistical and clustering techniques (e.g., PCA, invariant mining) to deep sequence and graph models such as DeepLog Du et al. (2017) and DeepTraLogZhang et al. (2022), which capture temporal and structural patterns in service calls. Metric based approaches rely on supervised and unsupervised learning over CPU, memory, I/O, and response-time series, ranging from simple MLP classifiersNobre et al. (2023) to NLP-based span analysis and distributed trace profiling. Recent hybrid methods integrate logs, traces, and metrics to exploit complementary views of system behaviorLi et al. (2024). This research explores anomaly detection using performance metrics.
Performance metrics based approaches monitor resource and latency indicators such as CPU, memory, disk I/O, network, and response time, and apply ML to detect deviations. Supervised models like simple Multi-Layer Perceptrons trained on labelled metric traces can accurately identify resource bottlenecks and failures. Alternatively, methods leveraging natural language processing on trace spans treat latency values as sequences, applying NLP techniques to spot performance regressions without prior system knowledge.
@@ 2.3 Root Cause Analysis
Modern deep anomaly detectors for microservices not only flag abnormal behaviors but increasingly aim to pinpoint why they occur, leveraging causal inference, graph reasoning, and multi-source fusion. Causal-discovery approaches such as CausalRCA Xin et al. (2023) and RUN Lin et al. (2024) employ gradient-based or neural Granger methods to learn directed graphs from metrics and time series, then infer fine-grained metric-level causes within faulty services. Graph-pruning and reinforcement-learning-based RCA (TraceDiagDing et al. (2023)) achieve interpretable, end-to-end localization on large-scale systems by learning to trim irrelevant dependencies before causal analysis. Although these methods embed some interpretability (e.g., attention weights, pruned graphs), only a few works explicitly adopt post-hoc XAI tools for RCA in microservices, e.g., yRCA Soldani et al. (2023). This paper explores the potential of anomaly localization through explainability using existing post-hoc explainability techniques such as SHapley Additive exPlanations (SHAP)Lundberg and Lee (2017b).
@@ 3.1.1 Testbed and Data Collection Environment
The RS-Anomic dataset was created using RobotShop Instana (2022), a microservice application that implements containerized orchestration and monitoring techniques. RobotShop is an open-source project with an e-commerce web application built using a microservices architecture. All services are deployed as Docker containers on a single server, communicating through a Docker bridge network. The application comprises 12 services, with their dependencies illustrated in Figure 1. This application was used in Harlicaj (2021) to evaluate their model. The testbed was run on a server with a 40-core CPU and 64 GB Memory running Ubuntu 20.04.5. We used Prometheus Rabenstein and Volz (2015) for data acquisition. The Prometheus server was configured to poll data from the microservices every 5 seconds. We instrumented the response times in microservice calls using Prometheus client libraries. The instrumented services are highlighted in Figure 1. Using Prometheus client libraries response time metrics were exposed from the microservices over HTTP endpoints. The cumulative summation of response time and the number of service calls that occurred between the data polling interval for each communication link was recorded. Container Advisor (cAdvisor) Google (2022) was utilized to gather runtime metrics. This tool offers measurements related to resource usage, such as CPU, memory, disk Input/Output (I/O), and network I/O of Docker containers. Additionally, data from cAdvisor was also polled from the Prometheus server. Data collection spanned over six days under normal conditions. For each type of anomalous behaviour, data was collected over 90 minutes by injecting faults into specific services to simulate anomalies. Table 1 shows the features available in the RS-Anomic dataset. To ensure RS-Anomic covers an actual e-commerce application behaviour, we simulated a time-varying load with more users in peak hours and fewer in off-peak hours. We used Locust Locust (2022), an open-source load testing tool, for load generation. The load generation can be done by creating test scenarios that simulate user behaviour and defining the number of virtual users and requests per second that should be generated during the test.
@@ 3.1.2 Anomalous Behaviours Captured in RS-Anomic
We observed that network packet loss, out-of-order packets, and high latency anomalies do not clearly distinguish between normal and anomalous data. Hence, we increased the strength of these anomalies. For example, we increased the latency from 200ms to 1000ms as a 200ms latency anomaly is more challenging to distinguish from normal data points.
3.2 Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) Model
Reproducing anomalies for various microservice applications is often impractical. Hence, the recent work has leveraged unsupervised approaches to develop anomaly detection models for microservices. Following them, we develop an unsupervised framework that uses the performance metrics of normally operating services for model training. Our model, Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD), employs an autoencoder architecture to learn a compact representation of normal behaviour and identify deviations as anomalies at test time.
We consider a microservice system comprising $n$ services, each characterized by $k$ performance metrics. Over $t$ time steps, we collect data representing the system’s behavior. GAL-MAD is structured as an autoencoder with an encoder-decoder architecture. The encoder utilizes two Graph Attention Network (GAT) layers Velickovic et al. (2018) followed by a bidirectional Long Short-Term Memory (Bi-LSTM) layer to capture spatial and temporal dependencies among services. The decoder mirrors the encoder’s architecture with a Bi-LSTM layer followed by two GAT layers applied in reverse order. GAL-MAD is trained on normal data to minimize the reconstruction loss using Mean Squared Error (MSE). The overall architecture of GAL-MAD is shown in Figure. 2.
Let $\mathbf{X}^{(i)}\in\mathbb{R}^{n\times k}$ denote the feature matrix at time step $i$, where each row corresponds to a microservice’s features and $\mathbf{A}\in\{0,1\}^{n\times n}$ represent the adjacency matrix encoding the static topology of microservice interactions. We stack the feature matrices over $t$ time steps to form a 3D tensor.
$\mathcal{X}=\left[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\ldots,\mathbf{X}^{(t)}\right]\in\mathbb{R}^{t\times n\times k}$ (1)
The encoder processes the input tensor $\mathcal{X}$ through two sequential GAT layers to learn latent representations, followed by a Bi-LSTM layer to capture temporal dependencies.
For each time step $i\in\{1,\ldots,t\}$, the first GAT layer computes the new node embeddings as follows.
$\mathbf{H}^{(1)}_{i}=\text{GAT}_{1}\left(\mathbf{X}^{(i)},\mathbf{A}\right)\in\mathbb{R}^{n\times d_{1}}$ (2)
, where $\text{GAT}_{1}$ denotes the first GAT layer and $d_{1}$ is the output feature dimension of the first GAT layer. After stacking over $t$ time steps, the output of the first GAT layer, $\mathcal{H}^{(1)}$, is
$\mathcal{H}^{(1)}=\left[\mathbf{H}^{(1)}_{1},\mathbf{H}^{(1)}_{2},\ldots,\mathbf{H}^{(1)}_{t}\right]\in\mathbb{R}^{t\times n\times d_{1}}$ (3)
Similarly, the second GAT layer, $\text{GAT}_{2}$, processes $\mathcal{H}^{(1)}$ to produce $\mathcal{H}^{(2)}$
$\mathcal{H}^{(2)}=\left[\mathbf{H}^{(2)}_{1},\mathbf{H}^{(2)}_{2},\ldots,\mathbf{H}^{(2)}_{t}\right]\in\mathbb{R}^{t\times n\times d_{2}}$ (4)
, where $d_{2}$ is the output feature dimension of the second GAT layer.
To capture temporal dependencies, we reshape $\mathcal{H}^{(2)}$ to a sequence suitable for the Bi-LSTM layer as follows.
$\mathcal{H}^{(2)}_{\text{reshaped}}\in\mathbb{R}^{n\times t\times d_{2}}$ (5)
The Bi-LSTM processes each service’s sequence over time to produce $\mathbf{z}_{j}$
$\mathbf{z}_{j}=\text{BiLSTM}\left(\mathcal{H}^{(2)}_{\text{reshaped}}[j,:,:]\right)\in\mathbb{R}^{d_{z}}$ (6)
, for $j=1,\ldots,n$, where $d_{z}$ is the dimension of the embedding vector for each service. After stacking the embeddings, we get the final embedding of the encoder $\mathbf{Z}=\left[\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{n}\right]^{\top}\in\mathbb{R}^{n\times d_{z}}$. Figure 3 elaborates on the encoder of the GAL-MAD.